OpenStudy (kakarot):

What is the value of d? 2d – 4 = 4 a.)d=1 b.)d=7 c.)d=4 d.)d=3

OpenStudy (mynameisnemo):

oh why did you close the question? @Kakarot do you still need help?

OpenStudy (kakarot):

why is this closed?

OpenStudy (kakarot):

I have no idea. I was lagging

OpenStudy (mynameisnemo):

I dont know but i can help you if you'd like

OpenStudy (kakarot):

ok.

OpenStudy (mynameisnemo):

Okay so this is how you would solve that: 2d - 4 = 4 + 4 + 4 2d = 8 /2 /2 d = 4

OpenStudy (mynameisnemo):

Understand how I got that answer?

OpenStudy (kakarot):

yes. 2 times 4 is 8. 8-4=4

OpenStudy (mynameisnemo):

are you proofing it or telling me thats how I did it?

OpenStudy (kakarot):

proofing it

OpenStudy (mynameisnemo):

Okay good XD